OpenStudy (anonymous):

Find the solution to the system of equations: 2x - 2y + z = -15 6x - 3y- z = -19 3x - y - z = -6

OpenStudy (anonymous):

X=-3 Y=2 Z=5

OpenStudy (anonymous):

just do the elimination 2x-2y+z=-15 2x-2y+z=-15 6x-3y-z=-19 3x-y-z = -6 ______________+ ______________+ 8x-5y=-34 5x-3y=-21 8x-5y=-34 >> 24x-15y=-102 5x-3y=-21 >> 25x-15=-105 ---------- ________________- -x=3 x=-3 and do the substitution 8(-3)-5y=-34 -24-5y=-34 -5y=-10 y=2 2(-3)-2(2)+z=-15 -6-4+z=-15 z=-5

OpenStudy (anonymous):

sorry for the typo on the first one the answer is x=-3 y=2 z=-5